OpenStudy (anonymous):

4. Find the volume of the cylinder in terms of . h = 6 and r = 3 (1 point) 27 in.3 108 in.3 54 in.3 324 in.3

OpenStudy (anonymous):

V=B*h, volume=base*height B is the area of the base which is a circle Area of a circle is pi*r^2 so take area of circle and multiply by height and that's the volume

OpenStudy (anonymous):

If that is done the that would give me 9pi because the formula is A= pi*r^2 witch is pi*3^2 witch then becomes 9pi.

OpenStudy (anonymous):

Yes, area of base is \[9\]

OpenStudy (anonymous):

9pi

OpenStudy (anonymous):

so then you would do 6*9 to get 54. so the answer would be 54in^3

OpenStudy (anonymous):

Not 54 but 54*pi

OpenStudy (anonymous):

oh okay now its showing me so it would be 54pi in ^3